# GDI-Bench: A Benchmark for General Document Intelligence with Vision and Reasoning Decoupling

### v1

@@ Abstract
The rapid advancement of multimodal large language models (MLLMs) has profoundly impacted the document domain, creating a wide array of application scenarios. This progress highlights the need for a comprehensive benchmark to evaluate these models’ capabilities across various document-specific tasks. However, existing benchmarks often fail to locate specific model weaknesses or guide systematic improvements. To bridge this gap, we introduce a General Document Intelligence Benchmark (GDI-Bench), featuring 1.9k images across 9 key scenarios and 19 document-specific tasks. By decoupling visual complexity and reasoning complexity, the GDI-Bench structures graded tasks that allow performance assessment by difficulty, aiding in model weakness identification and optimization guidance. We evaluate the GDI-Bench on various open-source and closed-source models, conducting decoupled analyses in the visual and reasoning domains. For instance, the GPT-4o model excels in reasoning tasks but exhibits limitations in visual capabilities. To address the diverse tasks and domains in the GDI-Bench, we propose a GDI Model that mitigates the issue of catastrophic forgetting during the supervised fine-tuning (SFT) process through a intelligence-preserving training strategy. Our model achieves state-of-the-art performance on previous benchmarks and the GDI-Bench. Both our benchmark and model will be open source.
@@ 1 Introduction
To address these challenges, we propose the General Document Intelligence Benchmark (GDI-Bench), which aims to perform a decoupled evaluation of model performance in document tasks, thereby contributing to the identification of the model’s weaknesses. The GDI-Bench introduces three key improvements over existing benchmarks: (1) developing a cross-domain, multi-task benchmarks to ensure task diversity and fine-grained difficulty levels; (2) introducing complexity decoupling, dividing multimodal document understanding into visual complexity and reasoning complexity, and for the first time, establishing a difficulty grading mechanism; (3) supporting the evaluation of MLLMs, OCR+LLM-level systems, and document parsing tools, offering comprehensive guidance for practical application solutions.
We evaluate 2 open-source and 4 closed-source large models using the GDI-Bench and find their performance to be suboptimal. To enhance the model’s performance on GDI-Bench, we further conduct supervised fine-tuning (SFT) on the InternVL3-8B model to explore whether a data-driven approach could help the model acquire extensive domain-specific document knowledge. However, we observe that supervised fine-tuning tends to cause the issue of catastrophic forgetting. To address this problem, we propose the Layer-wise Adaptive Freezing Tuning (LW-AFT) method, which alleviates the impact of catastrophic forgetting and enhances the model’s cross-domain and cross-task capabilities. Specifically, during the SFT process, LW-AFT freezes most of the parameters, with only a small subset of domain-sensitive parameters participating in gradient updates.
@@
  • We propose the GDI-Bench, which encompasses a broad range of tasks in the document domain. By decoupling complexity and grading difficulty, it helps the model identify its weaknesses and guides subsequent optimization.
@@
  • We propose a GDI model that achieves state-of-the-art (SOTA) performance on multiple document domain benchmarks as well as the GDI-Bench, demonstrating high generalization capabilities suitable for real-world applications.
@@ 2.1 Document Benchmark
Early benchmarks in document understanding predominantly targeted single-domain data and specific tasks. For instance, DocVQA[13] focused on industrial document QA, VisualMRC[14] addressed web-based document comprehension, and ChartQA[15] specialized in chart-based question answering. Despite their contributions, these benchmarks lacked difficulty stratification and cross-domain generalization, largely due to the limited generalization capabilities of models available at the time. With the rise of multimodal large models, a series of more complex benchmarks emerged [6, 7, 8, 9, 10, 11, 12], aiming to provide a more comprehensive framework for document understanding tasks. OCRBenchV2[16] provides multiple OCR-related subtasks, expanding the scope of OCR evaluation. Fox[17] introduced region-of-interest (ROI)-based document parsing methods, enhancing the detailed analysis of document structures, while OmniDocBench[18] extended the evaluation to cross-modal tasks, covering a broader range of document parsing challenges. While these benchmarks have broadened the scope of evaluation, they still lack explicit difficulty grading, which limits their ability to assess model performance across tasks of varying complexity.
2.2 Document Understanding Model
@@
Optical Character Recognition (OCR) has long been a fundamental task in computer vision. Existing OCR models can be broadly categorized into component-based and end-to-end approaches. Component-based methods [19, 20, 21] adopt a modular pipeline that assembles multiple expert-designed components such as layout analysis [22], text detection [23, 24, 25, 26], region extraction, and content recognition [27, 28, 29]. In contrast, end-to-end OCR models [30], especially those driven by Large Vision-Language Models (LVLMs)[31, 3, 32, 33, 34, 35, 36], aim to unify perception and reasoning within a single architecture. Most LVLM-based OCR systems utilize CLIP [37] as the vision backbone, while coupling it with a language model to jointly process visual and textual information in a unified framework. Recent works [34, 35, 38] adopt sliding-window strategies that partition the image into patches to cope with long and high-resolution inputs like PDFs.
2.3 Continual Learning of LLMs
@@
Continual learning in large language models (LLMs) faces the core challenge of catastrophic forgetting. Existing research mainly adopts two strategies: data replay  [39, 40] and parameter freezing [41, 42, 43]. Data replay strategies mitigate forgetting by revisiting samples from previous tasks during the training of new tasks. A typical approach is LAMOL[39], which generates pseudo-samples using the language models to avoid storage-based replay. Other approaches, such as experience replay and interleaved task training, have also been used to incorporate old data into the training loop. The parameter freezing method, on the other hand, protects existing knowledge by restricting parameter updates, such as through fine-tuning lightweight modules like LoRA[44], adapters [45], or prompt tokens [46]. Classical regularization-based methods like Elastic Weight Consolidation (EWC)  [47] protect crucial weight parameters through importance constraints. More recent methods, such as Task Vector [48] and Gradient Projection [49], operate on the gradient or representation level to minimize interference between tasks and preserve generalization. In contrast to these methods, which generally rely on a limited number of pre-training datasets, leading to incomplete domain coverage and significant computational costs, our method presents a more comprehensive and efficient alternative.
@@ 3.1 Complexity Decoupling
As illustrated in Figure 1, a difficulty-decoupled evaluation protocol named GDI-Benchmark is introduced to comprehensively assess MLLMs’ capabilities in visual information comprehension and reasoning. Distinct from VisualSimpleQA[50]’s structural decomposition of fact-seeking question answering tasks into visual recognition and knowledge dimensions through VLLM model architecture analysis, the proposed framework characterizes task complexity from fundamental difficulty perspectives by decoupling it into visual complexity and reasoning complexity components. Notably, knowledge popularity is explicitly subsumed within the reasoning complexity dimension in this formulation.
@@ 3.1.1 Visual Complexity
Figure 2: Distribution of visual complexity scores.
The visual complexity dimension is operationalized through a hierarchical categorization of document images into three levels: V0 (plain text), V1 (formal representations), and V2 (explanatory representations). V0 exclusively contains unstructured textual elements such as headings and paragraphs. Multimodal tasks on V0 documents typically achieve satisfactory performance via OCR-LLM pipeline architectures. As demonstrated in Figure 2, systematic analysis of the Omnidocbench benchmark reveals statistically significant performance gaps. This benchmark covers nine document categories. Current pipeline tools and VLLMs show notably worse performance on textbook, exam paper, notes, and newspaper types compared to other forms. These performance drops directly correlate with complex features like multi-column layouts and graphical elements. Based on these observations, we designate these four challenging categories as V2 and classify the remaining five as V1. This creates a data-driven taxonomy for visual complexity characterization.
@@ 3.1.2 Reasoning Complexity
The reasoning complexity characterization is formulated through a behavior-driven taxonomy. Three distinct levels are defined to progressively evaluate document understanding capabilities. It is specifically categorized into R0: Full Page Structured Extract, R1: Information Extract, and R2: Reasoning. Among these, the R0 task in V0 is more similar to OCR, and requires the ability to understand layout information such as tables, formulas, and complex page structures in V1 and V2 type images. For R1 tasks, the model needs not only to recognize visual content but also to understand the task itself in order to accurately extract relevant information, this often involves interpreting lines, tables, labels, and other visual elements present in V1 and V2 images. R2 tasks go a step further by requiring deeper reasoning capabilities, including the comprehension of logical inference and the ability to synthesize information across different modalities or layouts.
@@ 3.2 Annotation Process
3.2.1 Data Source
For the data acquisition phase of the General Document Intelligence (GDI) Benchmark, documents were primarily sourced from Omnidocbench[18], a comprehensive dataset containing document images from 9 domains. The dataset was further supplemented with an in-house collection of various document types, including exam papers, reports, newspapers and so on. This multi-source integration resulted in a well-rounded and representative dataset that captures the multifaceted nature of real-world document understanding tasks.
3.2.2 Data Construction
As shown in Figure 3, the data construction process begins with cropping single-layout sub-images from Omnidocbench[18] and inhouse documents to form the V0 raw image set. Based on end-to-end edit distance scores from state-of-the-art models and pipeline tools on Omnidocbench, domains with scores above 0.142 are categorized as V2, indicating high visual complexity. The remaining samples are assigned to V1.
For task construction, R0 tasks are generated using the original annotations or by synthesizing Markdown representations through MinerU[19]. To create R1 and R2 tasks, both the Markdown and corresponding images are input into GPT-4o to generate extractive and reasoning-based question-answer pairs. In addition, rule-based tasks are designed manually. These include extracting questions with fixed numbering, such as identifying problem numbers in exam papers.
To ensure high data quality, all synthetic cases are first evaluated by models. Cases that can be answered correctly by DeepSeek without requiring document input or that are flagged as low quality are filtered out. Finally, a team of PhD-level annotators reviews and verifies the remaining instances to ensure accuracy and correctness in the final benchmark dataset.
The final GDI-Benchmark contains a total of 2,989 test cases, distributed across different visual complexity levels (V0, V1, V2) and task types (R0, R1, R2), as summarized in Table 1.
Table 1: Complexity Distribution in the GDI Benchmark
@@ 3.2.2 Data Construction
r0 379 528 453 1 360
r1 223 407 276 906
r2 133 311 279 723
Total 735 1 246 1 008 2 989
3.3 Evaluation Metrics
Metrics are defined per task type. R0 uses Average Edit Distance (AED) for character-level recognition accuracy. R1 adopts Average Normalized Levenshtein Similarity (ANLS) for entity and field extraction evaluation. For R2, Accuracy is the primary metric, as most tasks require single-choice answers. A small number of structured reasoning tasks in R2 require models to generate structured outputs, for which AED is used to assess similarity to reference answers.
3.4 Comparison with Existing Document Intelligence Benchmarks
Table 2: Comparison of GDI-Bench with Other Document Understanding Benchmarks
@@ 3.4 Comparison with Existing Document Intelligence Benchmarks
GDI-Bench 9 19 1.9k

### v2

@@ Abstract
The rapid advancement of multimodal large language models (MLLMs) has profoundly impacted the document domain, creating a wide array of application scenarios. This progress highlights the need for a comprehensive benchmark to evaluate these models’ capabilities across various document-specific tasks. However, existing benchmarks often fail to locate specific model weaknesses or guide systematic improvements. To bridge this gap, we introduce a General Document Intelligence Benchmark (GDI-Bench), featuring 2.3k images across 9 key scenarios and 19 document-specific tasks. By decoupling visual complexity and reasoning complexity, the GDI-Bench structures graded tasks that allow performance assessment by difficulty, aiding in model weakness identification and optimization guidance. We evaluate various open-source and closed-source models on GDI-Bench, conducting decoupled analyses in the visual and reasoning domains, revealing their strengths and weaknesses. To address the diverse tasks and domains in the GDI-Bench, we propose a GDI-Model that mitigates catastrophic forgetting during the supervised fine-tuning (SFT) process through an intelligence-preserving training strategy, thereby reinforcing the inherent weaknesses of the base model. Our model achieves state-of-the-art performance on previous benchmarks and the GDI-Bench. Both our benchmark and models are or will be open-sourced on https://huggingface.co/GDIBench.
@@ 1 Introduction
To address these challenges, we propose the General Document Intelligence Benchmark (GDI-Bench), illustrated in Fig. 1, which aims to perform a decoupled evaluation of model performance on document tasks, thereby contributing to the identification of the model’s weaknesses. The GDI-Bench introduces three key improvements over existing benchmarks: (1) developing a cross-domain, multi-task benchmarks to ensure task diversity and fine-grained difficulty levels; (2) introducing complexity decoupling, dividing multimodal document understanding into visual complexity and reasoning complexity, and for the first time, establishing a difficulty grading mechanism; (3) supporting the evaluation of MLLMs, OCR+LLM-level systems, and document parsing tools, offering comprehensive guidance for practical application solutions.
Figure 1: Overview of GDI-Bench. The benchmark decouples document understanding complexity into visual complexity (V0-V2) and reasoning complexity (R0-R2) dimensions, creating a comprehensive evaluation framework for assessing MLLMs’ capabilities across various document types and reasoning tasks. Queries marked with a “CN” tag originate in Chinese and have been translated into English using Google Translate.
We evaluate 2 open-source and 4 closed-source large-scale models using the GDI-Bench and find their performance to be suboptimal, as shown in Fig. 2. For instance, while the InternVL3-8B model performs well in the R0 domain but shows significant performance degradation in the R1 and R2 domains To address the model’s weaknesses and further validate GDI-Bench’s capacity for weakness localization, we constructed supervised fine-tuning (SFT) on the InternVL3-8B model to explore whether data-driven approaches could improve its performance. However, we observe that supervised fine-tuning tends to cause the issue of catastrophic forgetting [13]. To address this problem, we propose the Layer-wise Adaptive Freezing-Tuning (LW-AFT) method, which alleviates the impact of catastrophic forgetting and enhances the model’s cross-domain and cross-task capabilities. Specifically, during the SFT process, LW-AFT freezes most of the parameters, with only a small subset of domain-sensitive parameters participating in gradient updates.
@@
  • We propose the GDI-Bench, a benchmark covering a broad range of document-related tasks. By decoupling complexity and grading difficulty, it helps the model identify its weaknesses and guides subsequent optimization.
@@
  • We propose a GDI-Model that achieves state-of-the-art (SOTA) performance on multiple document domain benchmarks as well as the GDI-Bench, demonstrating high generalization capabilities suitable for real-world applications.
Figure 2: Performance of various open-source and closed-source models on GDI-Bench at different levels of reasoning complexity. The GDI-Model is fine-tuned based on the InternVL3-8B model.
@@ 2 Related Works
Early benchmarks in document understanding primarily focused on single-domain tasks. For example, DocVQA[14] tackled industrial document QA, VisualMRC[15] focused on web-based comprehension, and ChartQA[16] specialized in chart-based QA. While these benchmarks made significant contributions, they lacked difficulty grading and cross-domain generalization, due to the limited capabilities of models at the time. With the advent of MLLMs, more complex benchmarks emerged [6, 7, 8, 9, 10, 11, 12], offering a more comprehensive framework for document understanding. OCRBenchV2[17] expanded OCR evaluation with various subtasks, while Fox[18] improved document parsing through ROI-based methods. OmniDocBench[19] extended evaluation to cross-modal tasks, covering a wider range of parsing challenges. Despite these advancements, these benchmarks still lack explicit difficulty grading, limiting their ability to assess performance across tasks of varying complexity.
Document Understanding Model
@@
Optical Character Recognition (OCR) has long been a fundamental task in computer vision. Existing OCR models can be broadly categorized into component-based and end-to-end approaches. Component-based methods [20, 21, 22] adopt a modular pipeline that assembles multiple expert-designed components such as layout analysis [23], text detection [24, 25, 26, 27], region extraction, and content recognition [28, 29, 30]. In contrast, end-to-end OCR models [31], especially those driven by MLLMs[32, 3, 33, 34, 35, 36, 37], aim to unify perception and reasoning within a single architecture. Most MLLM-based OCR systems utilize CLIP [38] as the vision backbone, while coupling it with a language model to jointly process visual and textual information in a unified framework. Recent works [35, 36, 39] adopt sliding-window strategies that partition the image into patches to cope with long and high-resolution inputs like PDFs.
Continual Learning of LLMs
@@
Continual learning in large language models (LLMs) faces the core challenge of catastrophic forgetting. Existing approaches mainly use two strategies: data replay  [40, 41] and parameter freezing [42, 43, 44]. Data replay mitigates forgetting by revisiting prior task samples during new task training. LAMOL[40], for example, generates pseudo-samples to avoid storage-based replay, while other methods like experience replay and interleaved task training integrate old data into the loop. Parameter freezing protects existing knowledge by limiting parameter updates, using techniques such as LoRA[45], adapters [46], or prompt tokens [47]. Regularization methods like Elastic Weight Consolidation (EWC) [13] preserve key weights via importance constraints, while newer methods like Task Vector [48] and Gradient Projection [49] operate at the gradient or representation level to reduce task interference and enhance generalization. In contrast to these approaches, which typically rely on limited pre-training datasets, our method offers a more efficient and comprehensive alternative.
@@ 3.1 Complexity Decoupling
As illustrated in Fig. 1, a difficulty-decoupled evaluation protocol named General Document Intelligence Benchmark (GDI-Bench) is introduced to comprehensively assess MLLMs’ capabilities in visual information comprehension and reasoning. Distinct from VisualSimpleQA[50]’s structural decomposition of fact-seeking question answering tasks into visual recognition and knowledge dimensions through MLLM model architecture analysis, the proposed framework characterizes task complexity from fundamental difficulty perspectives by decoupling it into visual complexity and reasoning complexity. Notably, knowledge popularity is explicitly subsumed within the reasoning complexity dimension in this formulation.
@@ 3.1.1 Visual Complexity
Figure 3: Distribution of visual complexity scores across nine document categories as reported in the OmniDocBench. The SOTA performance results from the benchmark show a substantial gap between document types, providing empirical justification for our visual complexity taxonomy.
The visual complexity dimension is operationalized through a hierarchical categorization of document images into three levels: V0 (plain text), V1 (formal representations), and V2 (explanatory representations). V0 exclusively contains unstructured textual elements such as headings and paragraphs. Multimodal tasks on V0 documents typically achieve satisfactory performance via OCR-LLM pipeline architectures. As demonstrated in Fig. 3, a systematic analysis of the OmniDocBench [19] benchmark reveals statistically significant performance gaps. This benchmark covers nine document categories. Current pipeline tools and MLLMs show notably worse performance on textbook, exam paper, notes, and newspaper types compared to other forms. These performance drops directly correlate with complex features like multi-column layouts and graphical elements. Based on these observations, we designate these four challenging categories as V2 and classify the remaining five as V1. This creates a data-driven taxonomy for visual complexity characterization.
@@ 3.1.2 Reasoning Complexity
The reasoning complexity characterization is formulated through a behavior-driven taxonomy. Three distinct levels are defined to progressively evaluate document understanding capabilities. It is specifically categorized into R0: Full Page Structured Extract, R1: Information Extract, and R2: Reasoning. Among these, the R0 task in V0 is more similar to OCR, and requires the ability to understand layout information such as tables, formulas, and complex page structures in V1 and V2 type images. For R1 tasks, the model needs not only to recognize visual content but also to understand the task itself in order to accurately extract relevant information. This often involves interpreting lines, tables, labels, and other visual elements present in V1 and V2 images. R2 tasks go a step further by requiring deeper reasoning capabilities, including the comprehension of logical inference and the ability to synthesize information across different modalities or layouts.
@@ 3.2 Annotation Process
3.2.1 Data Source
For the data collection phase of the GDI-Bench, documents are primarily sourced from OmniDocBench [19], a dataset containing document images from 9 domains. GDI-Bench is further supplemented with an in-house collection of various document types, including exam papers, reports, newspapers, and so on. This multi-source integration results in a well-rounded and representative dataset that captures the multifaceted nature of real-world document understanding tasks.
3.2.2 Data Construction
As shown in Fig. 4, the data construction process begins with cropping single-layout sub-images from OmniDocBench[19] and inhouse documents to form the V0 raw image set. Based on end-to-end edit distance scores from SOTA models and pipeline tools on OmniDocBench, domains with scores above 0.142 are categorized as V2, indicating high visual complexity. The remaining samples are assigned to V1.
@@ 3.2.2 Data Construction
For task construction, R0 tasks are generated using the original annotations or by synthesizing Markdown representations through MinerU [20]. To create R1 and R2 tasks, both the Markdown and corresponding images are input into GPT-4o to generate extractive and reasoning question-answer pairs. In addition, rule-based tasks are designed manually. These include extracting specific questions from exam papers, retrieving values enclosed in colored boxes, and so on.
To ensure high data quality, all synthetic cases are first evaluated by models. Cases that can be answered correctly by DeepSeek-R1[2] without requiring document input or that are flagged as low quality are filtered out. Finally, a team of PhD-level annotators reviews and verifies the remaining instances to ensure accuracy and correctness in the final benchmark.
The final GDI-Bench contains a total of 3,660 test cases, distributed across different visual complexity levels (V0, V1, V2) and task types (R0, R1, R2), as summarized in Table 1.
Table 1: Complexity Distribution in the GDI Benchmark.
@@
R0 377 521 450 1,348
R1 273 498 288 1,059
R2 298 641 314 1,253
Total 948 1,660 1,052 3,660
Table 2: Comparison of GDI-Bench with Other Document Understanding Benchmarks.
@@
GDIBench 9 19 2.3k
3.3 Evaluation Metrics
We adopt different evaluation metrics depending on the task type. For reasoning tasks, most questions are designed in a single-choice format to ensure standardized evaluation. For non-choice tasks, such as document parsing and extraction (mainly R0 and R1), where answers typically appear verbatim in the source document, we use the Normalized Edit Distance (NED) [51], but compute the score as $1-\text{NED}$ so that higher scores correspond to better performance. Formally, the case score is:
$\text{case}_{i}=\begin{cases}1-\text{NED}(\text{prediction}_{i},\text{ground truth}_{i})&\text{for non-choice tasks}\\ 1&\text{for choice-based tasks and correct answer}\\ 0&\text{for choice-based tasks and incorrect answer}\end{cases}$ (1)
The overall GDI-Bench Score is the average over all evaluation cases:
[…]
where $N$ is the total number of cases. A higher score indicates better overall performance.
3.4 Comparison with Existing Document Intelligence Benchmarks